Find an equivalent equation in rectangular coordinates. r=10sino
Answer #1

g=10sino ESAT Ince know narcoso ५ arsino n²+ y² = 2² 8=10 sino g²=10rsino 1²+ y² = 10 4 u+y²-10y =0 n²+ y² -10 4 20 is Equiva